Please answer part b! Thank you!
Phy284.Sp19 CS (p02 Correct 13
Answer #1

b) let one charge be q and other one be 4q

F = kq1q2 / r^2

0.2 = 9*10^9*q*4 / 0.11^2

q = 2.6*10^-7 C

Q1 = 2.6*10^-7 C

Q2 = 4q = 4*2.6*10^-7 = 1.04*10^-6 C